Can someone please explain all the steps of how to solve for (b) X=a+(b+2)w. I have seen several examples now but I still don't understand.
Answer by aditya2117(32)

X = a + (b+2)w
=> X-a = (b+2)w
=> X-a/w = b+2
=> b = X-a/w -2
therefore, b = [X-(a+2w)]/w

Now, putting the values of X,a,w, b is derrived.
For Example :
b = [10 - (5 + 2.1)]/1
= 10-7 = 3 [For X=10, a=5, w=1 ]
(ans.)